A coil of wire 8cm in diameter has 50turns and is placed in a B field of 1.8T. If the B field is reduced to 0.6T in 0.002seconds,what is the induced emf
1
Expert's answer
2020-05-20T09:08:53-0400
"\epsilon=NA\frac{B_1-B_2}{t}=0.25\pi d^2N\frac{B_1-B_2}{t}"

"\epsilon=0.25\pi (0.08)^2(50)\frac{1.8-0.6}{0.002}=150\ V"